OpenStudy (anonymous):

Write the quadratic equation in general quadratic form below. 3x2 + 6x = 12

OpenStudy (anonymous):

3x^2+6x-12=0 :D

OpenStudy (anonymous):

oh so its just ax^2+bx-c=d

OpenStudy (anonymous):

no it is \[ax^2+bx+c=0\] a zero on the right

OpenStudy (anonymous):

yea but she put -12 and 12 is c in this instance, right

OpenStudy (anonymous):

ax^2+bx+c=0..... c can be -12 or +12 or any constant value... here it turned out as -12

OpenStudy (anonymous):

right, but the point of subtracting 12 from both sides was to put a zero on the right